### v1

# Anomalous current fluctuations from Euler hydrodynamics

Abstract
We consider the hydrodynamic origin of anomalous current fluctuations in a family of stochastic charged cellular automata. Using ballistic macroscopic fluctuation theory, we study both typical and large fluctuations of the charge current and reproduce microscopic results which are available for the deterministic single-file limit of the models. Our results indicate that in general initial fluctuations propagated by Euler equations fully characterize both scales of anomalous fluctuations. For stochastic dynamics, we find an additional contribution to typical fluctuations and conjecture the functional form of the probability typical distribution, which we confirm by numerical simulations.
Introduction.—Statistical physics aims to describe universal behaviors emerging from interactions of a large number of microscopic degrees of freedom. While traditionally the main focus has been equilibrium physics, in recent years it has been appreciated that out-of-equilibrium properties of the system can also display rich universal behaviors. A prominent tool that describes such physics is macroscopic fluctuations theory (MFT) [1], which captures large fluctuations of driven diffusive systems. The validity of MFT has been verified for a variety of systems [2, 3, 4, 5], but large fluctuations in these systems usually behave in a regular way with finite scaled cumulants of observables. This naturally raises the question of whether there exists systems that break such regularity and support new dynamical phenomena.
Recent numerical studies have found dynamical criticality and anomalous spin fluctuations in integrable spin chains [6, 7], which have also been probed experimentally using superconducting qubits [8]. While the mechanism responsible for anomalous spin fluctuations remains to be clarified, an exact calculation of fluctuations in a deterministic charged cellular automaton [9] provided a basic example of dynamical criticality in a deterministic many-body system. The procedure has been generalized to a wider class of single-file dynamics, which was explored in Refs. [10, 11, 12, 13, 14, 15]. Furthermore, anomalous fluctuations have also been observed in the easy-axis regime of an integrable spin chain [7, 16] and in Dirac fluids [17], hinting at a more general setting supporting anomalous fluctuations.
Motivated by these, in this Letter, we provide a full hydrodynamic framework to understand the emergence of anomalous charge fluctuations in a family of stochastic charged cellular automata (SCCA), which was first introduced in Ref. [18] and reduces to the model studied in Ref. [9] in the single-file (deterministic) limit. The main tool for the analysis is ballistic MFT (BMFT) [19, 20], which is a generalization of MFT for systems that support ballistic transport, and we employ it to characterize both typical and large fluctuations. The simple three-mode hydrodynamic structure allows for an exact solution of BMFT equations, showing that large charge fluctuations are identical in the whole family of SCCA. While BMFT is in principle applicable only to large fluctuations, we also demonstrate that an approach based on the same idea, which is that current fluctuations are solely determined by initial fluctuations transported by Euler hydrodynamics, accurately describes typical fluctuations in the single-file limit, recovering the microscopic result [9]. Interestingly, unlike large fluctuations, it turns out that typical fluctuations depend on the stochasticity of dynamics. Based on numerical simulations, we conjecture a form of the corresponding typical distribution at half-filling.
Figure 1: (left) Stochastic particle scattering in the two-particle sector of the local two-body map $\Phi$. Particles with positive/negative charge (red/blue squares) either cross (top left) with probability $\Gamma$ or are elastically reflected (bottom left) with probability $\overline{\Gamma}$. Particle worldlines (colored lines) shown for clarity. (right) Many-body dynamics of charged particles and vacancies (black circles) in discrete space-time. Particles move along diagonals except when they encounter another particle when they scatter stochastically.
The model.—We consider a family of SCCA on a one-dimensional lattice of $L\in 2\mathbb{N}$ sites. The lattice configuration at fixed time $t\in\mathbb{Z}$ is given by strings ${\bf s}^{t}\equiv s^{t}_{L}s^{t}_{L-1}\ldots s^{t}_{1}$ of symbols $s^{t}_{x}\in\{\emptyset,-,+\}$, which correspond to vacancies and negatively or positively charged particles, respectively. Local symbol dynamics are given by a one-parameter stochastic two-body map $(s_{L}^{\prime},s_{R}^{\prime})=\Phi(s_{L},s_{R})$ which encodes the following dynamical rules (,)(,)($\emptyset$,$\emptyset$)$\rightarrow$($\emptyset$,$\emptyset$)( ∅ , ∅ ) → ( ∅ , ∅ ), $(\emptyset,c)\leftrightarrow(c,\emptyset)$ while $(c,c^{\prime})\to(c^{\prime},c)$ with probability $0\leq\Gamma\leq 1$ and $(c,c^{\prime})\to(c,c^{\prime})$ with probability $\overline{\Gamma}\equiv 1-\Gamma$ where $c,c^{\prime}\in\{-,+\}$, see Figure 1. We note that the dynamics can be understood as a composition of a stochastic six-vertex model in the particle-particle sector with free dynamics in the remaining sectors, see the Supplemental Material (SM) [21] for details.
The many-body dynamics is realized as a ‘brickwork’ circuit obtained by imposing periodic boundary conditions, $s^{t}_{x}=s^{t}_{L+x}$, and coupling alternating sites as ${\bf s}^{2t+2}=\Phi^{\rm odd}({\bf s}^{2t+1})$ and ${\bf s}^{2t+1}=\Phi^{\rm even}({\bf s}^{2t})$, where
$\Phi^{\rm odd}=\prod_{x=1}^{L/2}\Phi_{2x-1,2x},\qquad\Phi^{\rm even}=\prod_{x=1}^{L/2}\Phi_{2x,2x+1}$ (1)
and $\Phi_{x,x+1}$ acts non-trivially only on sites $x$ and $x+1$, see Figure 1 for a sample many-body trajectory. For $\Gamma=0$ and $\Gamma=1$ the dynamics is deterministic and corresponds to single-file [22, 23, 9, 10] and free dynamics respectively.
While the SCCA are super-integrable and support an exponential (in $L$) number of local conserved quantities [24], we presently restrict our study to a (closed) subset of conserved charges, comprising the number of right and left moving particles $\hat{Q}_{\pm}$ and the total charge $\hat{Q}_{c}$, all of which are given as sums of local densities $\hat{Q}_{i}({\bf s})=\sum_{x=1}^{L}\hat{q}^{i}_{x,t}({\bf s})$ where $i\in\mathcal{I}\equiv\{+,-,c\}$ and $\hat{\bullet}$ indicates a microscopic quantity, see SM [21] for details on charge densities and discrete space-time continuity relations.
We consider generalized Gibbs ensembles (GGEs) of the form $\mathbb{P}({\bf s})=\exp\left[\beta^{i}\hat{Q}_{i}({\bf s})\right]/\sum_{\bf s^{\prime}}\exp\left[\beta^{i}\hat{Q}_{i}({\bf s^{\prime}})\right],$ where repeated upper and lower indices indicate summation. The average of an observable $o({\bf s})$ is accordingly $\langle o\rangle=\sum_{{\bf s}}\mathbb{P}({\bf s})\,o({\bf s}).$ Identifying the ensemble parameters as $\rho_{\pm}=[1+e^{-\beta_{\pm}}/(2\cosh\beta_{c})]^{-1},b=\tanh\beta_{c}$, the ensemble factorizes $\mathbb{P}({\bf s})=\prod_{x=1}^{L/2}p_{-}(s_{2x+1})p_{+}(s_{2x})$ in terms of normalized one-site measures $p_{-}(\pm)=\rho_{-}\frac{1\pm b}{2}$, $p_{+}(\pm)=\rho_{+}\frac{1\pm b}{2}$, $p_{\pm}(\emptyset)=\overline{\rho}_{\pm}$ with $0\leq\rho_{\pm}\leq 1$ and $\overline{\rho}_{\pm}\equiv 1-\rho_{\pm}$ the densities of right/left movers and vacancies on right/left running diagonals, respectively.
Hydrodynamics.—The hydrodynamics of the system turns out to have a simple structure with densities of left and right moving particles $\varrho_{\pm}/2$ and charge $\varrho_{c}$ forming a closed set of hydrodynamic equations. To study charge transport (in the chosen GGE) it is then sufficient to consider the hydrodynamics of only these three charges unlike in a standard integrable systems where all charges must be included.
The three-mode hydrodynamic equations of the SCCA are obtained by evaluating the flux jacobian
$[\mathsf{A}[\underline{\rho}]]_{i}^{\leavevmode\nobreak\ j}=\frac{\partial j_{i}[\underline{\rho}]}{\partial\rho_{j}}=\begin{pmatrix}1&0&0\\ 0&-1&0\\ b\rho_{-}/\rho&-b\rho_{+}/\rho&v\end{pmatrix}$ (2)
where three modes are labeled by $\rho_{1}=\rho_{+}/2$, $\rho_{2}=\rho_{-}/2$, $\rho_{3}=\rho_{c}$ and the charge velocity $v$ is given by $v=p/\rho$ with $\rho=\rho_{1}+\rho_{2}$, $p=\rho_{1}-\rho_{2}$ the particle density and momentum respectively (see SM [21] for the derivation) and $j_{i}[\underline{\rho}]$ is the average current defined via the continuity equation $\partial_{t}\varrho_{i}+\partial_{x}j_{i}[\underline{\varrho}]=0$. Here, $\varrho_{i}$ is a space-time dependent density and the underline implies that the quantity is multicomponent. Charge inertness and chiral factorization of the dynamics amount to free ballistic propagation of the left and right movers $\partial_{t}(\varrho_{\pm}/2)\pm\partial_{x}(\varrho_{\pm}/2)=0$. On the other hand, charge transport is affected by particle dynamics and the Euler equation is of the form
$\partial_{t}\varrho_{c}+\partial_{x}(v\varrho_{c})=0.$ (3)
A system of three-mode hydrodynamic equations similar to Eq. (3) along with the free chiral equations for ϱ±subscriptitalic-$\varrho_{\pm}$ was also recently derived for Dirac fluids by keeping only the most relevant terms based on symmetry considerations in Ref. [17]. In the present case, the hydrodynamic equations are exact (i.e. no terms that contribute at the Euler scale are omitted), and interestingly, are completely insensitive to the crossing probability $\Gamma$. That said, as we will see later, typical fluctuations do turn out to depend nontrivially on $\Gamma$. Before discussing this dependence, we first study large fluctuations of the SCCA, which can be fully characterized using only Euler hydrodynamic data.
Anomalous charge FCS from BMFT.—The central object in the study of fluctuations is the generating function $\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle=\int{\rm d}\lambda\,e^{{\rm i}\lambda J}\mathcal{P}(J|t)$ given by the probability distribution $\mathcal{P}(J|t)$ of time-integrated charge current $\hat{J}(t)=\int_{0}^{t}{\rm d}t^{\prime}\,\hat{\jmath}_{c}(0,t^{\prime})$. While the scale on which typical fluctuations take place $\hat{J}(t)\sim t^{1/{2z}}$ is set by the variance $\langle[\hat{J}(t)]^{2}\rangle^{c}\sim t^{1/z}$, large (atypical) fluctuations occur on the scale $\hat{J}(t)\sim t$ with a probability distribution governed by a large deviation principle $\mathcal{P}({J}=jt|t)\asymp e^{-tI(j)}$ where $I(j)$ is the large deviation rate function. The rate function is related to the scaled cumulant generating function (SCGF)
$F(\lambda)=\lim_{t\to\infty}t^{-1}\log\langle e^{\lambda\hat{J}(t)}\rangle$ (4)
via the Legendre transform $I(j)=\max_{\lambda\in\mathbb{R}}[\lambda j-F(\lambda)]$. We now compute the charge SCGF in the SCCA explicitly using BMFT.
The fundamental postulate of BMFT is that large fluctuations of the initial state are propagated according to Euler equations in space-time, thereby governing the large-scale dynamics at later times. The idea is implemented by writing the generating function as
$\langle e^{\lambda\hat{J}(\tau)}\rangle=\int_{(\mathbb{S})}\mathcal{D}\underline{\varrho}(\cdot,\cdot)\mathcal{D}\underline{H}(\cdot,\cdot)e^{-\tau S[\underline{\varrho},\underline{H}]},$ (5)
where the functional integral is performed over the space of functions supported on the space-time region $\mathbb{S}=\mathbb{R}\times[0,1]$, and the densities $\varrho_{i}(x,t)$ are the fluctuating mesoscopic densities evaluated at the Euler-scale coordinates $(\tau x,\tau t)$, i.e. $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\tau x,\tau t)$. The auxiliary field $\underline{H}$ in Eq. (6) ensures that the fluctuating fields always satisfy the hydrodynamic equations, which in turn gives the BMFT action $S[\underline{\varrho},\underline{H}]$
$S[\underline{\varrho},\underline{H}]=-J(1)+\mathcal{F}[\underline{\varrho}(\cdot,0)]+\int_{\mathbb{S}}{\rm d}x{\rm d}t\,H^{i}(\partial_{t}\varrho_{i}+\partial_{x}j_{i}),$ (6)
where $J(1)=\int_{0}^{1}{\rm d}t\,j_{c}(0,t)$ is the time-integrated current and $\mathcal{F}[\underline{\varrho}(\cdot,0)]$ is the rate function of initial fluctuating densities
$\mathcal{F}[\underline{\varrho}(\cdot,0)]=\int_{\mathbb{R}}{\rm d}x\,\left((\beta^{i}(x,0)-\beta^{i})\varrho_{i}(x,0)-f+f(x,0)\right),$ (7)
with Lagrange multipliers $\beta^{i}(x,0)$, the free energy density $f(x,0)$ associated with the initial fluctuating density profile $\underline{\varrho}(\cdot,0)$, and the background free energy $f$. The saddle point of Eq. (5) now gives the BMFT equations
$\lambda\delta^{i}_{\leavevmode\nobreak\ c}\Theta(x)-\beta^{i}+\beta^{i}(x,0)-H^{i}(x,0)=0,$ (8)
$\lambda\delta^{i}_{\leavevmode\nobreak\ c}\Theta(x)-H^{i}(x,1)=0,$ (9)
$\partial_{t}\beta^{i}(x,t)+[\mathsf{A}^{\mathrm{T}}]^{i}_{\leavevmode\nobreak\ j}(x,t)\partial_{x}\beta^{j}(x,t)=0,$ (10)
$\partial_{t}H^{i}(x,t)+[\mathsf{A}^{\mathrm{T}}]^{i}_{\leavevmode\nobreak\ j}(x,t)\partial_{x}H^{j}(x,t)=0,$ (11)
where $\Theta(x)$ is the Heaviside step function. In terms of the solution $\underline{\beta}^{(\lambda)}(x,t)$ of the BMFT equations, the SCGF is computed as [19]
$F(\lambda)=\int_{0}^{\lambda}{\rm d}\lambda^{\prime}\int_{0}^{1}{\rm d}t\,j_{c}[\underline{\varrho}^{(\lambda^{\prime})}(0,t)]$ (12)
where $\underline{\varrho}^{(\lambda)}(x,t)=\underline{\varrho}[\underline{\beta}^{(\lambda)}(x,t)]$. Although solving BMFT equations is in general a formidable task even for (interacting) integrable systems, the simple hydrodynamic structure of the SCCA allows us to find their exact solution; see SM [21] for details. For $\beta^{\pm}=\beta$ and $t\in(0,1)$ the solutiong along the ray $x=0$ reads
$\beta^{(\lambda),c}(0,t)$ $=\beta^{c}(\lambda)=\beta^{c}+|\lambda|{\rm sgn}\,b,$ (13)
$\beta^{(\lambda),\pm}(0,t)$ $=\beta-2\Theta(\mp\lambda b)\log\left[\frac{\cosh\beta^{c}(\lambda)}{\cosh\beta^{c}}\right],$ (14)
which, using Eq. (12), also gives the SCGF
$F(\lambda)=\frac{1}{2}\log\left[1+\Delta^{2}(\mu_{b}+\mu_{b}^{-1}-2)\right]$ (15)
where $\Delta^{2}=\rho(1-\rho)$ and $\mu_{b}=\cosh\lambda+|b|\sinh|\lambda|$. This coincides precisely with the microscopic result obtained in [9] up to prefactor $1/2$ due to a different convention for the unit of time.
As was also observed in Ref. [9], the SCGF Eq. (15) is anomalous in that the scaled higher cumulants $c_{2n>4}=\lim_{t\to\infty}t^{-1}\langle[\hat{J}(t)]^{2n}\rangle^{c}$ are singular. BMFT clarifies the physical origin of this anomalous behavior: since the normal mode $b(x,t)$ of the charge $\varrho_{c}(x,t)$ travels along the ray $x=0$, fluctuations along $x=0$ proliferate and thereby render clustering of connected charge current correlations algebraic in time [25]. While such a scenario generally results in superdiffusion [26], it is avoided here due to the charge conjugation symmetry of the system [10]. In general, we expect that, for a charge in the system to show anomalous fluctuations, it has to have a vanishing velocity in equilibrium (usually afforded by a $\mathbb{Z}_{2}$ symmetry of the charge) and other modes should have no self-couplings, which would otherwise induce superdiffusion of the charge as in the heat mode with the $5/3$-Levy exponent in anharmonic chains [26]. It is also reasonable to assume that anomalous fluctuations are robust away from charge inertness, but a more precise analysis of such a situation is beyond the scope of the present Letter.
To our knowledge, the above solution is the first exact solution of full BMFT equations in an interacting model and therefore serves as an independent check of their validity. Furthermore, Eqs. (13) and (14) match the solution of the flow equation in ballistic fluctuation theory (see SM), corroborating the identification advocated in Ref. [19].
Note that Eq. (15) is valid for any value of $b\in(-1,1)$, including $b=0$, where the dynamical exponent becomes $z=2$. On the other hand, for typical fluctuations the two cases $b\neq 0$ and $b=0$ have to be treated separately since the dynamical exponents in the two cases differ. However, it turns out that, in the single-file (i.e., deterministic) limit $\Gamma=0$, the underlying mechanism, namely that initial fluctuations evolving according to the Euler equations fully determine charge current fluctuations, remains true even for typical fluctuations.
Figure 2: Finite-time typical distributions of the integrated charge current $t^{1/4}\mathcal{P}^{[\Gamma]}(Jt^{-1/4}|t)$ (colored lines) in linear and logarithmic scales at different $\Gamma$ compared against distributions $\mathcal{P}^{[\Gamma]}_{\rm typ}$ (20) (black lines) with $\omega$ and $\sigma$ as fitting parameters, best-fit values reported in the SM [21]. Simulation parameters: $\rho=1/2$, $b=0$, $L=2^{20}$, $t_{\rm max}=2^{16}$, $5\times 10^{3}$ samples.
Typical fluctuations of deterministic dynamics from Euler hydrodynamics.—The probability distribution of typical charge fluctuations $\hat{J}(t)\sim t^{1/2z}$ is given by
$\mathcal{P}(Jt^{-1/2z}|t)=\int_{\mathbb{R}}\frac{{\rm d}\lambda}{2\pi}\,e^{-{\rm i}\lambda Jt^{-1/2z}}\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle,$ (16)
which converges to $\mathcal{P}_{\mathrm{typ}}(j)=\lim_{t\to\infty}t^{1/2z}\mathcal{P}(J=jt^{1/2z}|t)$. We now evaluate Eq. (16) using only Euler hydrodynamic data, focusing on the single-file limit $\Gamma=0$, which has been discussed extensively in Refs. [9, 10, 11].
Recall that, the main ideas of both MFT and BMFT, which describe large fluctuations, hinge on the postulate that the generating function $\langle e^{{\rm i}\lambda\hat{J}(t)}\rangle$ can be expressed solely in terms of diffusive- or Euler-scale variables $\varrho_{i}(x,t)$, respectively, whose fluctuations are of order $O(1)$. Importantly, these fluctuating variables originate from atypical initial fluctuations described by Eq. (7), which evolve in time according to their respective hydrodynamic equations.
Motivated by this, we argue that, for large times, all contributions of the generating function to the integral in Eq. (16) for $\Gamma=0$ are captured by fluctuating variables $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\tau^{1/z}x,\tau t)$, $\tau\gg 1$, that can be obtained by propagating initial typical fluctuations:
$\varrho_{i}(x,0)\simeq\rho_{i}+\tau^{-1/2z}\delta\varrho_{i}(x,0),$ (17)
where $\rho_{i}$ is the initial average density and $\delta\varrho_{i}(x,0)$ is a field with $O(1)$ fluctuations.
The assertion that such fluctuations are always transported by the Euler equations regardless of the dynamical exponent stems from the observation that the only source of typical current fluctuations in the SCCA are the interactions of ballistically-propagating initial typical fluctuations. A similar observation was recently also made for Dirac fluids in Ref. [17]. This is in fact in the same spirit as diffusion from convection, which is known to be the only contribution to the diffusion constants in integrable systems [27, 28, 29] (we show that this is also the case in the SCCA in SM) and is generically also present even in non-integrable systems, provided that the eigenvalues of the flux jacobian $\mathsf{A}$ are non-degenerate [30].
To illustrate the idea, we focus on the case of half-filling $b=0$ and leave the details and $b\neq 0$ for the SM. Since in this case the dynamical exponent of charge transport is $z=2$, the fluctuating fields $\varrho_{i}(x,t)$ need to be rescaled to the diffusive coordinates $\varrho_{i}(x,t)=\hat{\varrho}_{i}(\sqrt{\tau}x,\tau t)$ with the associated currents $j_{i}(x,t)=\sqrt{\tau}\hat{\jmath}_{i}(\sqrt{\tau}x,\tau t)$. The fields $\varrho_{i}(x,t)$ then fluctuate initially as $\varrho_{i}(x,0)\simeq\rho_{i}+\tau^{-1/4}\delta\varrho_{i}(x,0)$, where $\delta\varrho_{i}(x,0)$ is Gaussian distributed as $\langle\delta\varrho_{i}(x,0)\delta\varrho_{j}(y,0)\rangle=\mathsf{C}_{ij}\delta(x-y)$ with the susceptibility matrix $\mathsf{C}_{ij}=\partial\rho_{i}/\partial\beta^{j}$ (see SM for the explicit expression). The fluctuating particle fields $\delta\varrho_{\pm}(x,0)$ are then transported by the Euler equation $\partial_{t}\delta\varrho_{\pm}(x,t)\pm\sqrt{\tau}\partial_{x}\delta\varrho_{\pm}(x,t)=0$, which can be solved as $\delta\varrho_{\pm}(x,t)\simeq\rho_{\pm}+\tau^{-1/4}\delta\varrho_{\pm}(x\mp\sqrt{\tau}t,0)$. Likewise the charge field $\delta\varrho_{c}(x,t)$ also fluctuates as $\delta\varrho_{c}(x,t)\simeq\tau^{-1/4}\delta\varrho_{c}(x-X_{d}(t),0)$ where $X_{d}(t)$ is the characteristics that satisfies ${\rm d}X_{d}(t)/{\rm d}t=\sqrt{\tau}v(X_{d}(t),t)$. With these we obtain the leading behavior of the fluctuating charge current at the origin $j_{c}(0,t)\simeq\tau^{-1/4}\partial_{t}\int_{0}^{X_{d}(t)}{\rm d}x\,\delta\varrho_{c}(-x,0)$, which in turn allows us to write down the functional integral that describes the contribution to the typical fluctuations by these fields
$\langle e^{\lambda\hat{J}_{c}(\tau)}\rangle$ $\simeq\frac{1}{Z}\int_{(\mathbb{R})}\mathcal{D}\underline{\delta\varrho}(\cdot,0)e^{-\frac{\mathsf{C}^{ij}}{2}\int_{\mathbb{R}}{\rm d}x\delta\varrho_{i}(x,0)\delta\varrho_{j}(x,0)}$
×$\quad\times e^{\lambda\tau^{1/4}\int_{0}^{X_{d}(1)}{\rm d}x\,\delta\varrho_{c}(-x,0)},$ (18)
where $\mathsf{C}^{ij}=[\mathsf{C}^{-1}]_{ij}$ and $Z$ is the normalization factor of the initial probability distribution. This integral can be performed exactly (see SM), yielding the non-Gaussian distribution
$\mathcal{P}_{\mathrm{typ}}(j)=\frac{1}{2\pi\Delta}\int_{\mathbb{R}}\frac{{\rm d}u}{\sqrt{|u|}}e^{-\frac{u^{2}}{2\Delta^{2}}-\frac{j^{2}}{2|u|}},$ (19)
which matches the result reported in Ref. [9]. The breakdown of Gaussianity in Eq. (19) can be understood by noting which fluctuating fields are responsible for charge current fluctuations. At half-filling, it follows from Eq. (18) that charge fluctuations $\delta\varrho_{c}(x,t)$ that travel along the characteristics $X_{d}(t)$ and cross $x=0$ between $t=0$ and $t=\tau$ contribute to current fluctuations. Since charge fluctuations are transported non-linearly, Gaussianity of inital fluctuations is no longer preserved and typical fluctuation become non-Gaussian. The situation changes dramatically away from half-filling, i.e. for $b\neq 0$ where only particle fluctuations $\delta\varrho_{\pm}(x,t)$ emanating along light cones contribute to charge current fluctuations. Since particle fluctuations are transported linearly, Gaussianity of initial fluctuations is preserved during the dynamics, giving rise to Gaussian typical fluctuations (see SM for details).
Typical fluctuations of stochastic dynamics.—The distribution (19) features a cusp at the origin. Numerical simulations for $\Gamma>0$ at half-filling indicate that the cusp is rounded off with stochastic crossings while the distribution remains non-Gaussian. This motivates a generalization of the single-file distribution (19) of the form
$\mathcal{P}^{[\Gamma]}_{\rm typ}(j)=\frac{1}{2\pi\sigma}\int_{\mathbb{R}}\frac{{\rm d}u}{\sqrt[4]{u^{2}+\omega^{2}}}e^{-\frac{u^{2}}{2\sigma^{2}}-\frac{j^{2}}{2\sqrt{u^{2}+\omega^{2}}}},$ (20)
which reduces to the single-file distribution (19) as $\omega\to 0$ and to a Gaussian distribution (with variance $\omega$) as $\omega\to\infty$. Treating $\omega$ and $\sigma$ as $\Gamma$-dependent fitting parameters, the distribution (20) accurately captures the asymptotic typical distribution for $\Gamma\geq 0$ across 5 order of magnitude as shown in Fig.  2.
Conclusions.—We have studied anomalous charge current fluctuations in a family of stochastic cellular automata and characterized both large and typical fluctuations based only on Euler hydrodynamics, reproducing the microscopic results obtained in Ref. [9] when specializing to the single-file (i.e. deterministic) limit $\Gamma=0$. The key insight is that initial fluctuations, carried by Euler hydrodynamics, fully control fluctuations both on the typical and large scales in the single-file case. To implement the idea, we employed BMFT to compute the SCGF and proposed a BMFT-based approach to describe typical fluctuations induced by initial Gaussian fluctuations (in a forthcoming work Ref. [31], it is also shown how BMFT provides the full diffusive-order hydrodynamic equation, where the long-range correlations predicted by BMFT are the source of diffusive-order terms). Our work clarifies for the first time the hydrodynamic origin of anomalous fluctuations, which we believe is generic among systems that exhibit anomalous fluctuations.
While initial fluctuations remain the only source of large fluctuations even in the stochastic case $\Gamma\neq 0$, it turns out that typical fluctuations are nontrivially influenced by the presence of stochasticity, which is not prescribed by initial fluctuations. We expect that stochasticity induces normal diffusion, and it would be very interesting to reproduce Eq. (20) by extending the formalism we introduced to incorporate normal diffusion. The observation of the distribution (19) in a classical spin chain [7] therefore supports the absence of normal diffusion in the easy-axis regime [32, 33]. Other important directions are to obtain the exact FCS of the SCCA via microscopic computations and establish a scattering description of thermodynamics, as recently obtained for classical integrable spin chains [34].
Acknowledgements.—We thank Bruno Bertini, Benjamin Doyon, Sarang Gopalakrishnan, Katja Klobas, and Tomaž Prosen for useful discussions. We are also grateful to Alvise Bastianello for his comments on the manuscript. ŽK is supported by the Simons Foundation as a Junior Fellow of the Simons Society of Fellows (1141511). This paper is dedicated to the memory of Marko Medenjak (1990-2022).